two charges of value 2mC and -50 mC are placed at a distance 80cm apart. . calculate the distance of the point from smaller charge where the intensity will be zero.

two  charges  of value 2mC  and -50 mC are placed at a distance 80cm apart. . calculate  the  distance of the point from smaller charge where the intensity  will be  zero.

Grade:12

1 Answers

Vikas TU
6 years ago
Amount of first charge is = 2* 10^-6 CAmount of second charge is = 50*10^-6 CDistance = 80 cm Electric Field = kq1q2/rElectric Field will be zeroE.F. = 0(K x 2* 10^-6 *50*10^-6)/80 + ((k*2* 10^-6)/r) = 0(10^-4/80) + (2x10^-6/r) = 010^-4/80 = -(2x10^-6/r)10^2/80 = 2/r10/8 = 2/r10/4 = 1/rR = 0.4